OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). y = one thirdx − 3 y = one thirdx + 17 y = −3x − 3 y = −3x + 17

OpenStudy (anonymous):

i really need help!

OpenStudy (anonymous):

Pleas help me

OpenStudy (loser66):

the given line is 3x +y =7---> y = -3x +7 tell me what is the slope of that line?

OpenStudy (anonymous):

6 right

OpenStudy (anonymous):

@loser66 is it 6?

OpenStudy (loser66):

I 'm sorry, I step away for awhile. nope, the slope of that line is -3, so the slope of the line perpendicular to that line is 1/3

OpenStudy (anonymous):

oh ok so you do the inverse @loser66

OpenStudy (loser66):

you have the form of the required line is y = 1/3 x + b. plug the point (6,-1) into that function to find b you have -1 = 1/3 *6 +b --> b = -3. then plug back to y , so you have y = 1/3 x -3 is solution

OpenStudy (anonymous):

@loser66 so the answer is the first choice?

OpenStudy (anonymous):

@loser66 hello?

OpenStudy (anonymous):

Yes it is the first choice

OpenStudy (anonymous):

@MDoodler thank u! could you also, help me on another question? i really dont understand parralel and perpendicular lines

OpenStudy (dan815):